• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Benaoum

    Log entry time 08:18:55 on September 2,2005

    Entry number 154214

    keyword=Owl Summary Shift



    03:09 Lost both arms and target due to network problem. Couldn't even connect online to reboot them. Have to wait while to be able to connect. Finally, I could connect to reboot IOC 13 for target and IOC 14 & 48. Both Left arm and target came up after this.
    Had some problem in recovering the right arm even when I rebooted IOC 16 (green button in the middle of the room).
    Paged Bob and asked to take data with only left arm.

    03:46 Rebooted again IOC 16 and right came back.
    06:45 MCC called to say they are taking beam away around 7:00am.




    Production Runs :
    -----------------
    4929 42.0k good pairs IHWP IN
    4930 45.9k good pairs IHWP OUT
    4931 14.4k good pairs IHWP OUT
    4932 46.8k good pairs IHWP OUT ( left arm only).
    4933 8.2k good pairs IHWP OUT ( not good for both arms).
    4934 41.5k good pairs IHWP OUT
    4935 42.0k good pairs IHWP OUT
    4936 14.1k good pairs IHWP OUT (ended this run b/c beam was taken away).

    Pedestal Runs :
    -----------------
    4937 Pedestal run.
    4938 Pedestal run.
    Pedestal runs will continue....
    Target warm up in progress.